A cash register contains only five dollar and ten dollar bills. It contains twice as many five's as ten's and the total amount of money in the cash register is 680 dollars. How many ten's are in the cash register?
Answer by haileytucki(390)

2X*5+X*10=680
Multiply 2X by 5 to get 10X.
10X+X*10=680
Multiply X by 10 to get 10X.
10X+10X=680
Since 10X and 10X are like terms, add 10X to 10X to get 20X.
20X=680
Divide each term in the equation by 20.
(20X)/(20)=(680)/(20)
Simplify the left-hand side of the equation by canceling the common factors.
X=(680)/(20)
Simplify the right-hand side of the equation by simplifying each term.
X=34 tens
68 fives

Prove:
34*10= 340
68*5= 340
340+340=680